OpenStudy (alexdiaz561):

A planet has a mass of about 6.4 x 1023 kg, and its moon has a mass of about 9.6 x 1015 kg. If the magnitude of the gravitational force between the two bodies is 4.6 x 1015 N, how far apart are the planet and its moon?

OpenStudy (nguyenvui):

The gravitation formula of Newton is \[F = G\frac{ m_{1} m_{2}}{ r^{2} }\] F = the force between two bodies. G = gravitational constant. = about 6.67 *10^(-11) m1 and m2 = are the masses of the bodies interacting. r = distance between two bodies. u can read full description here https://en.wikipedia.org/wiki/Gravity Now I will seek the distance \[r = \sqrt{\frac{ G m_{1}m_{2} }{ F }} = \sqrt{\frac{ 6.67* 10^{-11}*6.4*10^{23}*9.6*10^{15} }{ 4.6*10^{15} }}\] = ABOUT 9.4*106 meter (not quite exactly) u can find your answer! feel free to help!!^^